OpenStudy (anonymous):

A farmer is planning on planting 1,230 acres of corn and wheat, but wants to plant 35% more corn than wheat. How much of each should be planted?

OpenStudy (anonymous):

Solve the following equations for c and w.\[\left\{\frac{c-w}{w}=.35,w+c=1230\right\}\]\[\{w\to 523.4,c\to 706.6\}\]